OpenStudy (amber15):

SIMPLIFY 2 + 3/4 DIVIDED BY 5- 1/9 FIND THE QUOTE

OpenStudy (anonymous):

\[\huge \frac{2+\frac{3}{4}}{5-\frac{1}{9}}\]

OpenStudy (amber15):